problem 13 in Matter and Interactions 4th edition. Sorry. Elevator accelerating upward at rate 3.2m/s^2. block mass30 kg hangs from the ceiling, 68kg box hangs from it. What are the tenions in the upper and lower ropes? i thought it was m(a+g), but that's not working.
Answer #1

tension in lower wire

T = m ( g + a)

T = 68 ( 9.8 + 3.2) = 884 N

======

tension in upper wire

t = (68 + 30)* ( 9.8 + 3.2) = 1274 N

======

Comment before rate in case any doubt, will reply for sure.. goodluck